# Beyond Formula Complexity: Effective Information Criterion Improves Performance and Interpretability for Symbolic Regression

### v1

@@ Abstract
Symbolic regression discovers accurate and interpretable formulas to describe given data, thereby providing scientific insights for domain experts and promoting scientific discovery. However, existing symbolic regression methods often use complexity metrics as a proxy for interoperability, which only considers the size of the formula but ignores its internal mathematical structure. Therefore, while they can discover formulas with compact forms, the discovered formulas often have structures that are difficult to analyze or interpret mathematically. In this work, inspired by the observation that physical formulas are typically numerically stable under limited calculation precision, we propose the Effective Information Criterion (EIC). It treats formulas as information processing systems with specific internal structures and identifies the unreasonable structure in them by the loss of significant digits or the amplification of rounding noise as data flows through the system. We find that this criterion reveals the gap between the structural rationality of models discovered by existing symbolic regression algorithms and real-world physical formulas. Combining EIC with various search-based symbolic regression algorithms improves their performance on the Pareto frontier and reduces the irrational structure in the results. Combining EIC with generative-based algorithms reduces the number of samples required for pre-training, improving sample efficiency by $2\sim 4$ times. Finally, for different formulas with similar accuracy and complexity, EIC shows a $70.2\%$ agreement with 108 human experts’ preferences for formula interpretability, demonstrating that EIC, by measuring the unreasonable structures in formulas, actually reflects the formula’s interpretability.
@@ 1 Introduction
Figure 1: Formulas with the same complexity but different interpretability. Formulas with similar complexity and accuracy can have very different structures and interoperability.
@@
Symbolic regression (SR) is a machine learning technique that discovers interpretable mathematical formulas describing relationships in data (camps2023discovering, ). Unlike black-box models, it reveals how inputs map to outputs in a form that can be analyzed mathematically or logically, offering insights and new scientific knowledge (makke2024interpretable, ). In a typical workflow, researchers apply SR algorithms to data to generate candidate formulas that balance accuracy and complexity (i.e., the Pareto front). They can then assess the credibility of these candidates and select the most interpretable formula for insights into the underlying patterns (liu2024automated, ). Traditionally, SR relies on heuristic search, especially genetic programming, which edits formulas through mutation and crossover to maximize a search objective that balances complexity and accuracy (cranmer2023interpretable, ; burlacu2020operon, ). Recently, pretraining-based methods have gained attention, in which transformer models are pre-trained on synthetic formula-data pairs to predict formula symbols from data. After training, these models can generalize to real systems, generating formulas directly (biggio2021neural, ; kamienny2022end, ; meidani2023snip, ) or guiding symbolic search (kamienny2023deep, ; shojaee2023transformer, ; yu2025symbolic, ).
Most SR methods use formula length, measured by the number of symbols in a formula (la2021contemporary, ; aldeia2025call, ), as a proxy for interpretability. However, the assumption that shorter formulas are more interpretable is not always valid. The length metric only captures overall size but ignores the internal structure of a formula, which strongly affects interpretability. For example, although the two formulas in Figure 1 have the same length and similar accuracy, the left one is difficult to interprete with its nested structure $\sin(\sin(\cot(x)))$, whereas the right one, as a linear combination of simpler functions, is more likely to provide scientific insights. Thus, while search-based algorithms can discover short formulas, they cannot ensure interpretable structures, limiting the process of extracting knowledge and hindering SR in scientific discovery. On the other hand, generative methods face a similar issue: pre-trained on random formulas with inevitably unreasonable structures, they typically show poor sample efficiency and require tens or hundreds of millions of samples to achieve effective generalization (kamienny2022end, ; yu2025symbolic, ).
To address this problem, we propose the Effective Information Criterion (EIC), grounded in information theory and numerical analysis, to identify unreasonable structures in formulas. The key idea is to view a formula as an information-processing system and evaluate its structural plausibility based on the loss of effective information it outputs under finite processing precision, relative to the maximum information a system could output under the same precision. This approach is motivated by the observation that human-derived physical equations can typically be computed with limited precision (e.g., manual calculations or floating-point arithmetic) without substantial loss of significant digits. In contrast, formulas learned by SR often exhibit unreasonable structures such as catastrophic cancellation, leading to a large loss of significant digits under limited precision. As shown in Figure 2, while real physical formulas typically lose less than one significant digit, SR results, despite being compact and accurate, can lose three or more, revealing prevalent structural issues. Building on this, we integrate EIC into multiple stages of the SR workflow. For heuristic algorithms, it serves as an auxiliary search objective to guide the search toward more structurally reasonable results. For generative methods, EIC can filter out formulas with unlikely structures during pre-training, reducing the number of samples needed for effective generalization and thus improving sample efficiency.
@@
Figure 2: EIC distributions of real physical formulas and existing SR results. The figure compares the EIC distribution of 133 ground-truth physical formulas with that of formulas discovered by existing SR algorithms. While real physical formulas typically have a low EIC of $0\sim 1$, formulas discovered by SR often exceed 3, highlighting the prevalence of unreasonable structures. The inner box shows the quartiles, and the black dots indicate outliers beyond $1.5\timesIQR$.
We conduct extensive experiments to demonstrate EIC’s effectiveness in enhancing both the performance and interpretability of formulas learned by SR. First, unlike traditional complexity measures, EIC evaluates the plausibility of a formula’s internal structure, providing a more accurate assessment of interpretability and potential for scientific insight. Second, integrating EIC into heuristic search algorithms enhances search performance, yielding formulas that occupy superior positions on the Pareto front with fewer unreasonable structures. Third, filtering high-EIC samples during pre-training for generative methods significantly reduces the number of training samples required, improving sample efficiency by 2 to 4 times. Models trained on low-EIC samples also generalize better to real-world physical formulas, achieving an $R^{2}$ improvement of 22.4% compared to those trained on random formulas. Finally, in an evaluation of 172 formula pairs with similar $R^{2}$ and complexity but different EIC values, assessed by 108 human experts, we find that EIC aligns with human preferences $70.2\%$ of the time. This demonstrates that EIC can not only identify unreasonable structures but also evaluate interpretability effectively. By incorporating EIC into the search, training, and evaluation stages of SR, we can enhance both algorithm performance and the potential to extract scientific knowledge from learned formulas.
@@ 2 Related Work
Symbolic regression (SR) is a machine learning task that discovers symbolic formulas from data by combining variables, constants, and mathematical operators. Existing SR algorithms can be broadly divided into heuristic search-based and generative methods.
Heuristic search-based symbolic regression methods. Traditionally, SR is implemented using heuristic search methods such as genetic programming (augusto2000symbolic, ; schmidt2010age, ; de2021interaction, ; arnaldo2014multiple, ; burlacu2020operon, ; virgolin2019linear, ; kartelj2023rilsrols, ; smits2005pareto, ; la2016epsilon, ; cranmer2023interpretable, ; zhong2018multifactorial, ; burlacu2020operon, ; searson2010gptips, ; virgolin2021improving, ; zhang2022ps, ), Monte Carlo tree search (sun2022symbolic, ), and deep reinforcement learning (petersen2021deep, ; tenachi2023deep, ). These methods search for functions that balance simplicity and accuracy, producing candidate formulas for experts’ analysis. Some methods leverage neural networks to identify symmetries (udrescu2020ai, ), fit Q-functions (xu2023rsrm, ), or predict minimum description length (yu2025symbolic, ) to guide the search. While these methods often limit the maximal formula length to avoid overly complex formulas, they cannot prevent formulas with unreasonable structures, such as deeply nested nonlinear functions or catastrophic cancellation (i.e., subtracting nearly equal subformulas). Although there are methods using unit constraints to ensure physically reasonable structures (tenachi2023deep, ), they are largely limited to physics. In broader domains like ecology, sociology, and earth science, where variable units may be unclear and models can include many constants with unknown units, such constraints are difficult to apply. This motivates incorporating EIC into the search objective to exclude unreasonably structured formulas, improving the interpretability of discovered results.
@@
Generative symbolic regression methods. In recent years, generative methods have emerged, which pre-train transformers on large-scale randomly generated formula-data pairs to predict symbols in underlying formulas given data. Pre-trained models can generate formulas (biggio2021neural, ; kamienny2022end, ), guide symbolic search (shojaee2023transformer, ; kamienny2023deep, ; yu2025symbolic, ; ying2025neural, ; yu2025neural, ), or serve as foundation models for downstream tasks (meidani2023snip, ). Despite their promise, these methods face significant challenges in sample efficiency. Randomly generated formulas often contain unreasonable structures that differ from real-world physical formulas, requiring extremely large training datasets (tens or hundreds of millions) for effective generalization. This motivates using EIC to filter out formulas with unreasonable structures during pre-training, thereby improving consistency with real-world formulas and enhancing sample efficiency.
3 Methodology
In this section, we introduce the proposed Effective Information Criterion (EIC), illustrated in Figure 3. Section 3.1, corresponding to the upper part of the figure, introduces its definition and physical meanings from perspectives of numerical analysis and information-processing systems. Section 3.2, corresponding to the lower part, introduces three ways EIC can enhance symbolic regression workflows and facilitate scientific discovery.
Figure 3: Demonstration of the physical meanings of EIC and the way it can enhance symbolic regression workflows.
3.1 Effective Information Criterion (EIC)
Definition. We view a symbolic formula not merely as a static mathematical object, but as an information-processing system. A formula maps inputs to outputs through a sequence of mathematical operations and thus forms an information flow. Under limited processing precision, rounding errors inevitably occur at operator nodes and propagate through the computation graph. In well-structured formulas, these errors remain controlled, so the effective information at the output approaches the system’s maximum. In unreasonably structured formulas, however, small intermediate perturbations can accumulate and severely degrade the output’s effective information. This motivates evaluating a formula’s structure by how stably it processes information and how much effective information is lost under finite precision.
To quantify this, we introduce the Effective Information Criterion (EIC), which measures information loss in terms of significant digits. If a formula computed with $N$ significant digits produces an output with only $M$ significant digits, its EIC is defined as
$\mathrm{EIC}\triangleq N-M.$ (1)
Therefore, a large EIC indicates a formula that amplifies rounding errors and leads to severe information loss, while a small EIC indicates that the formula structure preserves most of the output information. For example, $f(x)=(x+10^{100})-10^{100}$ suffers catastrophic cancellation when computed with fewer than $100$ significant digits, leaving no reliable information output. In contrast, physical formulas from science typically preserve output information with finite precision, resulting in small EIC values.
Calculation. Limited processing precision can be modeled as noise injected into intermediate computations. Specifically, truncating a variable $x$ to $N$ significant digits can be expressed as $\tilde{x}=x+e$, where the error term $e$ behaves like uniform noise following $\mathcal{U}(-5\times 10^{M-N},5\times 10^{M-N})$, with $M=\log_{10}|x|$ representing the order of magnitude of $x$. Expressed as relative noise $e/x$, it intensity is independent of $x$ (and $M$):
$\sigma_{r}^{2}\triangleq\mathrm{Var}\left[\frac{e}{x}\right]=\frac{1}{12}10^{2(1-N)},$ (2)
or equivalently,
$N=1-\frac{1}{2}\log_{10}(12\sigma_{r}^{2}).$ (3)
This shows that retaining a certain number of significant digits can be converted into injecting noise of a specific intensity. This equivalence provides us a practical method for calculating EIC: we first compute the formula outputs $\tilde{y}$ with noise of relative strength $\sigma_{r}^{2}$ added. Then, we estimate the relative noise strength at the output through $\delta_{r}^{2}\triangleq\mathrm{Var}\left[\frac{\tilde{y}-y}{y}\right]$, where $y$ is the noiseless output. Finally, the output noise can be converted back to significant digits as $M=1-\frac{1}{2}\log_{10}(12\delta_{r}^{2})$, using the relationship above.
Physical meanings. The correspondence between significant digits and relative noise intensity provides both a method for calculating EIC and a richer physical interpretation. Specifically, from equation 3, we have (see Appendix A.1 for details)
$\mathrm{EIC}=N-M=\log_{10}(\delta_{r}^{2}/\sigma_{r}^{2}).$ (4)
This highlights EIC’s dual meaning: from a numerical analysis perspective, it measures the number of significant digits lost (i.e., NMN-M), while from an information-processing perspective, it quantifies the relative noise gain along the formula’s computational graph (i.e., $\log_{10}(\delta_{r}^{2}/\sigma_{r}^{2})$). In other words, EIC captures the sensitivity of a formula’s output to small intermediate perturbations, thus identifying formulas that may be compact but mathematically fragile. Both perspectives emphasize EIC’s value as a physically meaningful criterion for detecting unreasonable mathematical structures.
Implementaion details. EIC can be calculated recursively. As shown in Appendix Algorithm 1, computing a formula’s EIC requires computing the noisy and noise-free outputs of each subformula first, which further depend on their subformulas, down to the input variables and constants. This allows EIC to be computed for both the final output and each subformula. We found that in some cases, subformulas have higher EICs than the full formula. For example, $f(x)=0\times(x+10^{100}-10^{100})+x$ contains catastrophic cancellation inside the parentheses, but the zero coefficient nullifies its effect, resulting in a small EIC at the output. To account for such cases, we use the maximum EIC among all subformulas as the formula’s overall EIC.
3.2 Enhance Symbolic Regression with EIC
Improve heuristic search-based methods. EIC evaluates unreasonable structures in formulas and can be integrated as an auxiliary objective in classical heuristic search algorithms to guide the search and improve performance. We focus on two representative approaches: genetic programming (GP) and Monte Carlo tree search (MCTS), which underpin many state-of-the-art methods such as PySR(cranmer2023interpretable, ), TPSR(shojaee2023transformer, ), SR4MDL(yu2025symbolic, ), etc. GP and MCTS iteratively edit formulas to maximize a search objective, balancing accuracy and complexity. A commonly used fitness function is defined as
$f(C,\text{NMSE};\eta)={\eta^{C}}/{(1+\text{NMSE})},$ (5)
where $C$ denotes the formula complexity (length), $\text{NMSE}=\text{MSE}/\text{Var}(y)$, and $\eta<1$ penalizes formula complexity (sun2022symbolic, ) (see Appendix A.2 for details). It is worth noting that, when calculating the MSE, we first decompose formulas into additive terms and then fit their coefficients using linear regression, avoiding more costly nonlinear optimization algorithms such as BFGS(fletcher2000practical, ) and accelerating the search. EIC can be incorporated by modifying the search objectives:
$f(C,\text{NMSE},\text{EIC};\eta,\alpha)=\eta^{\text{Complexity}}/(1+\text{NMSE})-\alpha\cdot\text{EIC},$ (6)
where $\alpha>0$ penalizes formulas with higher EIC, steering the search away from structurally unreasonable solutions.
Improve sample efficiency of generative methods. Generative approaches, such as NeuralSR(biggio2021neural, ), E2ESR (kamienny2022end, ), and SNIP (meidani2023snip, ), are typically pretrained on synthetic formula-data pairs, aiming to predict tokens (i.e., symbols) in formulas from input data. These methods often rely on the random formula generation algorithm developed by lampledeep , which produces diverse symbolic expressions for pretraining. While this enables generating unlimited training data for pre-training, many synthetic formulas contain mathematically unreasonable structures rarely seen in real-world physical equations. Therefore, these models often require tens to hundreds of millions of samples and weeks of computation for effective generalization(kamienny2022end, ; meidani2023snip, ). To address this, we propose filtering the pretraining corpus via rejection sampling based on EIC. By discarding formulas with high EIC values, we remove structurally implausible samples while preserving overall diversity. Specifically, during pretraining, we evaluate the EIC value of each generated sample. Samples exceeding a specified threshold are regenerated until the EIC is below the threshold.

### v2

# Beyond Accuracy and Complexity: The Effective Information Criterion for Structurally Stable Symbolic Regression

@@ Abstract
Symbolic regression (SR) traditionally balances accuracy and complexity, implicitly assuming that simpler formulas are structurally more rational. We argue that this assumption is insufficient: existing algorithms often exploit this metric to discover accurate and compact but structurally irrational formulas that are numerically ill-conditioned and physically inexplicable. Inspired by the structural stability of real physical laws, we propose the Effective Information Criterion (EIC) to quantify formula rationality. EIC models formulas as information channels and measures the amplification of inherent rounding noise during recursive calculation, effectively distinguishing physically plausible structures from pathological ones without relying on ground truth. Our analysis reveals a stark structural stability gap between human-derived equations and SR-discovered results. By integrating EIC into SR workflows, we provide explicit structural guidance: for heuristic search, EIC steers algorithms toward stable regions to yield superior Pareto frontiers; for generative models, EIC-based filtering improves pre-training sample efficiency by 2–4 times and boosts generalization $R^{2}$ by 22.4%. Finally, an extensive study with 108 human experts shows that EIC aligns with human preferences in 70% of cases, validating structural stability as a critical prerequisite for human-perceived interpretability.
@@ 1 Introduction
Symbolic regression (SR) is a machine learning technique that discovers interpretable mathematical formulas describing relationships in data (Camps-Valls et al., 2023). Unlike black-box models, it reveals how inputs map to outputs in a form that can be analyzed mathematically or logically, offering insights and new scientific knowledge (Makke and Chawla, 2024). In a typical workflow, researchers apply SR algorithms to data to generate candidate formulas that balance accuracy and complexity (i.e., the Pareto front). They can then assess the credibility of these candidates and select the most reasonable formula for insights into the underlying patterns (Liu et al., 2024). Traditionally, SR relies on heuristic search, especially genetic programming, which edits formulas through mutation and crossover to maximize a search objective that balances complexity and accuracy (Cranmer, 2023; Burlacu et al., 2020). Recently, generative methods have gained attention, in which transformer models are pre-trained on synthetic formula–data pairs to predict formula symbols from data. (Biggio et al., 2021; Kamienny et al., 2022; Meidani et al., 2023). However, trained on synthetic formulas with inevitably irrational structures, these methods typically struggle with sample efficiency and require tens or hundreds of millions of samples to achieve effective generalization in real-world systems (Kamienny et al., 2022; Yu et al., 2025). Recent approaches shift back to the search-based paradigm, integrating pre-trained neural networks (Kamienny et al., 2023; Shojaee et al., 2023; Yu et al., 2025), online-trained deep reinforcement learning policies (Petersen et al., 2021; Tenachi et al., 2023; Xu et al., 2024), or large language models (Shojaee et al., ) to guide the search algorithms.
These search-based SR methods typically evaluate candidate formulas based on the trade-off between accuracy and complexity (i.e., the Pareto frontier) (La Cava et al., 2021). However, this standard bi-objective framework relies on the implicit assumption that lower complexity implies higher rationality, which is not always valid. The complexity metric, usually measured by the number of symbols in a formula (i.e., formula length) (La Cava et al., 2021; Aldeia et al., 2025), captures only the overall size but ignores the internal structure of a formula, which is critical for physical plausibility. Driven by powerful search capabilities, existing SR algorithms often discover accurate and compact, but structurally irrational formulas. As illustrated in Figure 1, although the two formulas possess identical complexity and similar accuracy, they exhibit drastically different structural characteristics: the left one features a nested, irrational structure (e.g., $\sin(\sin(\cot(x)))$) that is physically inexplicable and numerically ill-conditioned, whereas the right one, as a linear combination of simpler functions, aligns well with scientific intuition. The prevalence of such structurally irrational results in practice limits the process of extracting knowledge and hinders the application of SR in scientific discovery.
Figure 1: Formulas with identical complexity and accuracy can exhibit distinct structural rationality. Despite having the same length and $R^{2}$, the left formula contains pathological nesting, while the right one remains structurally sound, demonstrating that complexity fails to capture structural rationality.
To address this problem, we draw inspiration from the structural stability of real physical formulas. In Figure 2, we observe a stark structural stability gap: while human-derived equations are typically robust to computational noise (e.g., floating-point errors), formulas discovered by SR, despite achieving high accuracy and compactness, often suffer from critical structural instability. These mathematically pathological structures drastically amplify rounding noise and, critically, render the formulas irrational and uninterpretable. Consequently, we propose the Effective Information Criterion (EIC) to quantify structural stability. EIC measures the amplification of cumulative noise generated when inherent rounding noise in the calculation propagates in the formula, thereby effectively distinguishing physically plausible formulas from irrational ones.
@@
Figure 2: The structural stability gap between real physical formulas and SR results. We compare 133 ground truth formulas (Feynman, Strogatz) with results from 17 SR algorithms, quantified by the noise amplification factor (Section 3). The inner box shows the quartiles, and the black dots indicate outliers beyond $1.5\timesIQR$.
EIC reveals a stark structural stability gap between real physical formulas and those discovered by existing SR methods. This finding motivates us to integrate EIC into SR workflows to mitigate structural irrationality. Specifically, for heuristic search-based algorithms, EIC guides the search toward stable regions, yielding superior Pareto frontiers while effectively mitigating structural instability. For generative methods, it filters out unstable samples during pre-training, boosting sample efficiency by 2–4 times and enhancing generalization to physical formulas by an $R^{2}$ of 22.4%. Finally, an evaluation involving 108 human experts confirms that EIC aligns with human preferences in 70% of cases, validating that structural stability is a critical prerequisite for human-perceived interpretability.
@@ 2 Related Work
Existing symbolic regression (SR) algorithms can be broadly divided into heuristic search-based and generative methods. Additionally, their metrics and constraints are critical to understanding current limitations.
Heuristic search-based SR methods. Traditionally, SR relies on heuristic search methods such as genetic programming (Augusto and Barbosa, 2000; Burlacu et al., 2020; Cranmer, 2023) or Monte Carlo tree search (Sun et al., 2022). Recently, advanced approaches have integrated pre-trained neural networks (Kamienny et al., 2023; Shojaee et al., 2023; Yu et al., 2025; Ying et al., 2025), online-trained deep reinforcement learning policies (Petersen et al., 2021; Tenachi et al., 2023; Xu et al., 2024), or large language models (Shojaee et al., ) to guide the search algorithms. Furthermore, some works leverage neural networks to identify symmetries in data or perform ab initio variable selection (Ye and Li, 2024) to facilitate the search process (Udrescu and Tegmark, 2020). While these methods typically limit the maximal formula length to avoid excessive complexity, they can barely prevent formulas with irrational structures. This motivates us to incorporate the proposed EIC into the search objective to penalize structurally unstable candidates, thereby improving the rationality of the discovered results.
@@
Generative symbolic regression methods. In recent years, generative methods have emerged, which pre-train Transformer models on large-scale randomly generated formula – data pairs to predict mathematical symbols directly from data. These pre-trained models can serve various roles, including generating formulas directly (Biggio et al., 2021; Kamienny et al., 2022), guiding symbolic search algorithms (Shojaee et al., 2023; Kamienny et al., 2023; Yu et al., 2025; Ying et al., 2025), or acting as foundation models for downstream tasks (Meidani et al., 2023). Despite their promise, these methods face significant challenges in sample efficiency. The randomly generated training formulas often contain irrational structures that diverge significantly from real-world physical laws, thus requiring extremely large datasets (tens or hundreds of millions of samples) to achieve effective generalization. This motivates us to use EIC to filter out formulas with irrational structures during the pre-training phase, thereby enhancing consistency with real-world formulas and boosting sample efficiency.
Evaluation metrics and constraints used in SR. Standard SR benchmarks, such as SRBench (La Cava et al., 2021), SRBench++ (de Franca et al., 2024), SRSD (Matsubara et al., 2022), and LLMSRBench (Shojaee et al., 2025), typically employ a bi-objective framework balancing fitting accuracy (e.g., $R^{2}$, RMSE) and complexity (Fong and Motani, 2024). Common complexity metrics like formula length, depth, and minimum description length (Yu et al., 2025) capture overall size but, as Figure 1 illustrates, ignore internal structure. Moreover, standard evaluations of out-of-distribution (OOD) performance or noise sensitivity focus solely on input-output mappings, overlooking internal rationality. For instance, while $x$ and (10000x9999x)(10000x-9999x) yield identical mappings, the latter suffers from catastrophic cancellation, amplifying rounding noise by 4 orders of magnitude. Finally, regarding physical plausibility, methods imposing unit constraints (Tenachi et al., 2023) are largely limited to physics and difficult to apply in domains like sociology, where units are ambiguous. Meanwhile, manual expert evaluations of trustworthiness either reflect subjective preferences (La Cava et al., 2022) or are too time-consuming for standard benchmarks (Virgolin et al., 2023). Consequently, no scalable metric currently evaluates internal structural rationality. EIC fills this gap by explicitly measuring structural stability, complementing existing complexity-based and black-box evaluations.
3 Effective Information Criterion
3.1 Intuition and Definition for EIC
Motivation. As white-box models, symbolic formulas are defined critically by their internal structure, which determines their rationality (Figure 1), rather than just input-output mappings. We draw inspiration from history: foundational laws were verified using low-precision tools (e.g., slide rules with 3–4 significant digits), implying a strong inductive bias that valid formulas must remain robust under limited precision. Consequently, we propose evaluating rationality via structural stability. Viewing a formula as a computation tree (with inputs as leaves and operators as nodes), intrinsic rounding errors inevitably accumulate and amplify during propagation. Unlike real physical formulas, which are typically stable, SR-discovered models, even those with accurate fitting, often possess ill-conditioned structures that amplify computational noise by orders of magnitude (Figure 2), which renders them irrational. Therefore, we propose quantifying this noise amplification to effectively distinguish real physical formulas from irrational ones.
Mathematical Formalization. Formally, we represent a formula $f$ as a symbol tree $\mathcal{T}[f]$ (as demonstrated in Figure 1). Each internal node $k$ implements an elementary operator $e_{k}$ (e.g., ++, $\times$, $\sin$) with child nodes $\mathcal{C}[k]$. Given an input sample $x$ drawn from a data distribution $\mathcal{D}$, the ideal, noise-free value $y_{k}(x)$ at node $k$ can be obtained recursively via $y_{k}(x)=e_{k}(\{y_{i}(x)\}_{i\in\mathcal{C}[k]})$, where $e_{k}(\cdot)$ represents the ideal operator output computed under infinite precision. Under finite precision, however, the computed value $\tilde{y}_{k}$ at node $k$ deviates from $y_{k}$ due to the inherent rounding noise introduced at this node and the propagated error accumulated from the child nodes. Since the magnitude of rounding error is proportional to the value itself, it can be modeled as a multiplicative noise with zero mean and a fixed variance:
Assumption 3.1 (Rounding noise as multiplicative noise).
$\tilde{y}_{k}=e(\{\tilde{y}_{i}\}_{i\in\mathcal{C}[k]})\times(1+\epsilon_{k}),$ (1)
where $\epsilon_{k}$ is an independent random variable [^1] Although rounding error is deterministic for a fixed input $x$, quantization theory suggests that as computational precision increases (i.e., variance of the rounding error decreases), the error behaves pseudo-randomly and becomes asymptotically uncorrelated with $x$ (Widrow et al., 1996) following any distribution with $\mathbb{E}[\epsilon_{k}]=0$ and $\text{Var}[\epsilon_{k}]=\sigma^{2}$.
To quantify the extent to which the introduced rounding error is amplified due to accumulation during propagation, we define the cumulated error $\eta_{k}\triangleq(\tilde{y}_{k}-y_{k})/y_{k}$ and the variance amplification factor
$s^{2}_{k}(x)\triangleq\lim_{\sigma\to 0}\frac{\text{Var}[\eta_{k}|x]}{\sigma^{2}},$ (2)
which quantifies the sensitivity of the subtree rooted at node $k$ to rounding noise at high precision limits ($\sigma\to 0$) by the degree to which accumulated noise is amplified compared to rounding noise. Note that $s_{k}(x)$ is a function of $x$ since the same subtree structure can have different sensitivities under different input data. Intuitively, a small $s_{k}(x)$ suggests the subtree is structurally stable, while a large $s_{k}(x)$ suggests the subtree acts as a noise amplifier, significantly magnifying small perturbations from upstream calculations.
To analyze how the error accumulation is intrinsically driven by the mathematical structure of the formula, we derive its recursive calculation rule (See Appendix A.1 for detailed derivation):
Proposition 3.2 (Recursive Relation of $s_{k}(x)$).
$s_{k}^{2}(x)=1+\sum_{i\in\mathcal{C}[k]}\kappa_{k,i}^{2}(x)\times s_{i}^{2}(x),$ (3)
where $\kappa_{k,i}\triangleq\frac{y_{i}}{y_{k}}\frac{\partial e_{k}}{\partial y_{i}}$ is the partial relative condition number of the operation $e_{k}$ with respect to its operand $i$. For leaf nodes where $\mathcal{C}[k]=\emptyset$, we assume its $s_{k}^{2}(x)=1+0=1$.
This reveals that $s_{k}(x)$ is a deterministic structural property, computable solely from the formula’s structure and input data $x$ without relying on external parameters. Physically, it decomposes the total instability into two components, including the intrinsic rounding noise generated at the current node (represented by the constant term $1$) and the propagated noise from child nodes ($s_{i}^{2}(x)$), which is amplified by the squared condition number $\kappa_{k,i}^{2}$.
Definition of EIC. Based on the node-wide structural instability measure $s_{k}(x)$, we define the Effective Information Criterion (EIC) as
$\text{EIC}\triangleq\max_{k\in\mathcal{T}[f]}\log_{10}\bar{s}_{k}.$ (4)
Here, $\bar{s}_{k}^{2}\triangleq\mathbb{E}_{x\sim\mathcal{D}}[s_{k}^{2}(x)]$ averages local point-wise instability into a global metric. This expectation ensures the evaluation is contextualized within the scientifically relevant domain $\mathcal{D}$, accounting for nonlinear structures where stability fluctuates across regions (e.g., $1/x$ is more unstable when $x$ is smaller). The $\log{10}$ term maps $s_{k}\in[1,\infty)$ to a linear scale of precision loss, providing clear physical meaning. Crucially, we employ $\max_{k}$ rather than the root metric $s_{\text{root}}$ to impose a strict worst-case constraint. This prevents internal instabilities from being masked by insensitive downstream operations (specifically where $\kappa_{k,i}\approx 0$; see Appendix A.2). This definition, combined with Eq (3), allows recursively computing $s_{k}(x)$ and EIC with $O(N)$ linear complexity via post-order traversal (Appendix Algorithm 1).
Intuitive understanding of EIC. The EIC quantifies the structural stability of a formula $f$ over the data distribution $\mathcal{D}$ by the amplification of cumulative noise $\eta_{k}$ generated when inherent rounding noise $\eta_{k}$ in the calculation propagates in the formula. This can be seen more clearly through (see Appendix A.3 for proof)
$\mathrm{EIC}=\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}\log_{10}\frac{\mathrm{Std}[\eta_{k}]}{\mathrm{Std}[\epsilon_{k}]},$ (5)
where $\mathrm{Std}[\cdot]=\sqrt{\mathrm{Var}[\cdot]}$ is the standard deviation of random variables that reflects their strength. This relationship clearly illustrates that EIC is the magnitude by which the intensity of the accumulated noise is amplified compared to the added perturbation when a sufficiently small perturbation is added during the calculation of the formula, which is just as shown by the coordinate axes on both sides of Figure 2.
Theoretical Properties. Besides its intuitive understanding, EIC exhibits several desirable properties. First, it is parameter-free. As Eq. (3) shows, calculating $s_{k}(x)$ depends solely on structure and partial condition numbers, independent of external hyperparameters like noise level $\sigma$. Second, EIC is scale-invariant. Since condition numbers are dimensionless, EIC remains invariant under unit changes (e.g., rescaling meters to centimeters), as detailed in Appendix A.4. Finally, EIC is sensitive to algebraic rewriting. For instance, $f(x)=x$ and its redundant equivalent $g(x)=\ln(\exp(x))$ yield distinct values. Specifically, when $\mathbb{E}_{\mathcal{D}}[x]\triangleq\mu\gg 1$, we find $\text{EIC}(g)\approx\log_{10}(\mu)>\text{EIC}(f)=0$ (see Appendix A.5). Although mathematically equivalent, they differ as calculation graphs: $g$ introduces unnecessary operations that amplify noise. This aligns with parsimony but relies on computational stability rather than symbol counting.
3.2 Physical Interpretations
While the definition of EIC clearly elucidates its capability to quantify structural stability, linking it to established physical concepts can further strengthen its theoretical foundation
Numerical Precision Perspective. EIC measures structural stability under multiplicative noise, which can be physically interpreted as the loss of significant digits. According to quantization theory (Widrow et al., 1996), the operation of rounding to $N$ significant digits is statistically equivalent (in the variance sense) to injecting multiplicative noise with variance $\sigma^{2}\propto 10^{-2N}$. Leveraging this equivalence, we can formally link EIC to the reduction in effective precision (see Appendix A.6 for the detailed derivation):
$\text{EIC}=\lim_{N\to\infty}(N-\min_{k\in\mathcal{T}}M_{k}),$ (6)
where $N$ denotes the working precision in significant digits (corresponding to $\sigma^{2}$), and $M_{k}$ represents the effective significant digits remaining at node $k$ after noise accumulation. Therefore, EIC quantifies the maximum magnitude of the intrinsic precision loss imposed by the formula’s structure in the asymptotic limit of infinite precision ($N\to\infty$). As demonstrated in Figure 2, EIC values of SR-discovered formulas can reach up to 6688, implying a catastrophic intrinsic loss of 6688 significant digits. In sharp contrast, ground-truth physical laws typically exhibit $\text{EIC}<1$, reflecting a natural preference for structurally robust representations that preserve information fidelity.
Signal Processing Perspective. Complementing the static numerical precision view, a dynamic perspective treats the formula as a signal transmission channel, where the input data represents the source signal and the rounding errors act as channel noise. In this signal processing system, noise figure serves as a fundamental metric that quantifies the degradation of the Signal-to-Noise Ratio (SNR) as information traverses the system: $F_{k}\triangleq 10\log_{10}\frac{\text{SNR}_{\text{intrinsic}}}{\text{SNR}_{k}}$ (in decibels), where $\text{SNR}_{\text{intrinsic}}\triangleq 1/\sigma^{2}$ represents the theoretical upper bound of signal quality determined by machine precision, and $\text{SNR}_{k}\triangleq 1/\text{Var}[\eta_{k}]$ denotes the actual SNR remaining at node $k$ after structural noise amplification. As derived in Appendix A.6, EIC can be formally interpreted as the noise figure (scaled by $\frac{1}{20}$) at the bottleneck node:
$\text{EIC}=\frac{1}{20}\lim_{\sigma\to 0}\max_{k\in\mathcal{T}[f]}F_{k}.$ (7)
This mapping reveals a profound insight: a high EIC indicates that the formula acts as a noisy channel that severely corrupts the information flow. For instance, while real physical formulas act as high-fidelity channels (EIC <1<1, or $20\text{dB}$ loss) that preserve information integrity, a SR-discovered formula with a EIC $\approx 6$ acts as destructive channels with a $120\text{dB}$ SNR drop, implying that the signal quality (SNR) has degraded by twelve orders of magnitude (a $10^{12}\times$ loss in power) relative to the input and rendering the output signal virtually indistinguishable from noise.